Free cash flow valuation Nabor Industries is considering going public but is unsure of a fair offering price for the company.All parts on this question need answers.
Answer #1

a: Value of business= FCF1/(1+r)^1 + FCF2/(1+r)^2 …………FCFn/(1+r)^n+ Horizon value/(1+r)^n

= 240000/1.12^1+270000/1.12^2+320000/1.12^3+390000/1.12^4+430000/1.12^5+ (430000*103%/(12%-3%))/1.12^5

=3941513.94

b: Common stock value = Value of business- Debt-Preferred stock

= 3941513.94-1970000-390000

= 1581513.94

C:Stock price = 1581513.94/200000 = $7.91